OpenStudy (anonymous):

Radon-222 is a radioactive gas with a half-life of 3.82 days. How long would it take for fifteen-sixteenths of a sample of radon-222 to decay? 15.28 days 30.56 days 45.84 days 49.66 days

sam (.sam.):

\[\Huge t=\frac{\log\frac{A_E}{A_O}}{\log0.5} \times t_{\frac{1}{2}}\] t is the time it would take t1/2 is the half life AO is the amount of substance at orignial time AE is the amount of substanse left \[\Huge t=\frac{\log\frac{\frac{1}{16}}{1}}{\log0.5} \times3.82days\] \[\Huge t=15.28\]

OpenStudy (anonymous):

Thank you, that's the formula I needed.